OpenStudy (anonymous):

Consider the function f(x) = x^{2}e^{3 x}. For this function there are three important intervals: (-\infty, A], [A,B], and [B,\infty) where A and B are the critical numbers. Find A and B For each of the following intervals, tell whether f(x) is increasing (type in INC) or decreasing (type in DEC). (-\infty, A]: [A,B]: [B,\infty):

OpenStudy (dumbcow):

A function is Increasing when the slope is positive or f'(x) >0 A function is Decreasing when the slope is negative or f'(x) <0 The critical points occur when slope is zero ... f'(x) = 0 so step 1 : take derivative of function

OpenStudy (dumbcow):

use product rule: \[(fg)' = f'g + fg'\] where f = x^2 and g = e^3x

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

So what did you get for the derivative?

OpenStudy (anonymous):

f'(x)=2xe^3x+x^2e^3x(3)

OpenStudy (anonymous):

Yes... now factor out the greatest common factor in each term and set the whole thing equal to zero.

OpenStudy (anonymous):

e^3x(2x+3x^2)=0

OpenStudy (anonymous):

x comes out too

OpenStudy (anonymous):

\[e^{3x}x(2+3x)=0\]

OpenStudy (anonymous):

e^3x will never equal zero... but that x right next to it will... so xe^(3x)=0 and 2+x=0 solve for x in each and you will have A and B

OpenStudy (anonymous):

I dont know what I am doing

OpenStudy (anonymous):

You seem to know a bit... work from there. correction above... you are right... the term on the right it (2+3x), not 2+x. See... you know better than me! :)

OpenStudy (anonymous):

haha I know how to do the derivative thats about it

OpenStudy (anonymous):

and factor apparently :)

OpenStudy (anonymous):

Ohh i find the critical numbers

OpenStudy (anonymous):

OKay let me see.................

OpenStudy (anonymous):

so when two things are multiplied and equal to zero, either factor can make the whole thing zero.... yes this is how you find the crits

OpenStudy (anonymous):

Soi s one number 0

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The other comes from 2+3x=0, because e^(3x) can never be zero.

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

woo hoo

OpenStudy (anonymous):

woof!

OpenStudy (anonymous):

Ha okay I dont know what to do next

OpenStudy (anonymous):

What is A and B

OpenStudy (anonymous):

Cause I put in 0 and -2/3 and it said it was wrong

OpenStudy (anonymous):

Now I draw a number line with -2/3 and 0 on it. Find a test point in each interval that they specify and see if the derivative is positive or negative in those intervals... I used -1, -1/3, and 1. It is -2/3 first... then 0 for your program.

OpenStudy (anonymous):

ooh heh

OpenStudy (anonymous):

\[(-\infty, - \frac{2}{3}) U (- \frac{2}{3}, 0) U ( 0, \infty)\]

OpenStudy (anonymous):

Oh yes I just test those points to see right?

OpenStudy (anonymous):

I get it!!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Don't test the crits... you'll get zero... pick something to the sides of them that are easy to calculate.

OpenStudy (anonymous):

Yes, I knw wher to test

OpenStudy (anonymous):

Cool!

OpenStudy (anonymous):

Its INC DEC INC

OpenStudy (anonymous):

That's what I got! :)

OpenStudy (anonymous):



OpenStudy (anonymous):

woof! :)

OpenStudy (anonymous):

Can you help me with something else?

OpenStudy (anonymous):

Post a new question... then I'll look at it.

OpenStudy (anonymous):

okay